The first month of the winter in Bloomington, is a cold month, with an average temperature ranging between max. 5.3 °C and min. -4.1 °C. During this time there are a lot of snows and the roads are covered with ice. Which chemical compound(s) you use to lower the freezing point of ice for the streets of Bloomington and why? Suppose, you dissolve 87.0 g of rock salt in 500.0 g of water. What will be the freezing point of that solution?  (Kf = 1.86 oC/molal for water).


1
Expert's answer
2021-01-18T01:57:33-0500

"to - tf = \frac{w}{m}\times\frac{1000}{Wg}\times Kf"

"0 - tf = \frac{87}{58.5}\times\frac{1000}{500}\times 1.86"

"tf = -5.53"0C